OpenStudy (sparklestaraa):

In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2^x)

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (danjs):

You can fill in the table right?

OpenStudy (sparklestaraa):

uhhh do i just replace the like 0 with the x and all?

OpenStudy (danjs):

Yeah just put x= 0,1,2,3,4,5 into both and record the values

OpenStudy (sparklestaraa):

can u wait while i try it??

OpenStudy (sparklestaraa):

is 1.2^0 1?

OpenStudy (danjs):

Here is the graph for both the linear s(x), and the exponential m(x), functions.

OpenStudy (sparklestaraa):

:D

OpenStudy (sparklestaraa):

I can't do the exponent on the table can u do that for me ?

OpenStudy (danjs):

s(x) = 100 + 23x ----------------- Points (x , s(x)) (0 , 100) (1, 123) (2, 146) (3, 169) (4, 192) (5, 215)

OpenStudy (sparklestaraa):

this is b?

OpenStudy (sparklestaraa):

Thats confusing like where r we

OpenStudy (danjs):

m(x) = 90*(1.2)^x ------------------------ points (x, m(x)) (0, 90) (1, 108) (2, 129.6) (3, 155.5) (4, 186.6) (5, 223.9)

OpenStudy (danjs):

Those are all the values for the table.

OpenStudy (sparklestaraa):

wait so what?????

OpenStudy (sparklestaraa):

so the table is what? because i need to have 1 and 2 to the table

OpenStudy (sparklestaraa):

ahh i think ik

OpenStudy (danjs):

For x=0 to x=5, those are all the values for both of the functions...

OpenStudy (sparklestaraa):

i got it :P I understood halfway through my desperate typing xD

OpenStudy (danjs):

Part b, Using the table, where do both functions have the same value... It is between x=4, and x=5. So approx... when x=4.5 both functions will have the same value.

OpenStudy (danjs):

Because, m(x) goes from being smaller than s(x) , to larger than s(x) between x=4 and x=5, so somewhere in that interval, both functions must intersect and have the same value.

OpenStudy (danjs):

Part c, the graph is linked, but i think they want you to do it yourself on paper, you have to plot each of those points from the table, then draw a line and curve manually.

OpenStudy (danjs):

Here is the graph, you have to label each of the points from the table on the line and curve.

OpenStudy (sparklestaraa):

ohhhh so how do i do a curve? and lane EVERY point on what?

OpenStudy (sparklestaraa):

and *plot

OpenStudy (danjs):

Draw a graph like the one i linked, label every point from the table on the graph.

OpenStudy (danjs):

like for s(x) , the points are... (0 , 100) (1, 123) (2, 146) (3, 169) (4, 192) (5, 215)

OpenStudy (danjs):

for m(x), the points are... (0, 90) (1, 108) (2, 129.6) (3, 155.5) (4, 186.6) (5, 223.9)

OpenStudy (sparklestaraa):

ohhhh whats m(x) is that the curve?

OpenStudy (danjs):

The next part d) From the problem, the X axis values represent the Week #, week 0 (initial value) through week 5 The Y value represents the number of organisms present.

OpenStudy (sparklestaraa):

oh ok :P and b i think its just the intersection

OpenStudy (danjs):

Yes m(x) is an exponential equation

OpenStudy (sparklestaraa):

oh ok

OpenStudy (danjs):

Each point on the graph (x,y) = (week #, population of organisms)

OpenStudy (sparklestaraa):

this is still d ryt?

OpenStudy (danjs):

Part e, to determine the solution from the graph, you estimate the intersection point (x,y)

OpenStudy (danjs):

The last part, you actually have to do the algebra and solve for the exact value for the intersection point... \[100+23x=90(1.2)^x\]

OpenStudy (sparklestaraa):

OMG how on earth do i do that :O

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (sparklestaraa):

wait so is the exact value the values in the table?

OpenStudy (danjs):

The table are exact values for each function. the exact value for m(x) = s(x) is not on the table. (4.433, 201.97) is the decimal approx value for the answer.

OpenStudy (sparklestaraa):

ohhhh so how do u figure exact because i can guarantee u don't know xD

OpenStudy (danjs):

divide both sides by 90 take natural log of both sides ...

OpenStudy (sparklestaraa):

whats a log?

OpenStudy (sparklestaraa):

like natural log xD not a log

OpenStudy (danjs):

log base e

OpenStudy (danjs):

or you can use log base 10, it doesnt matter

OpenStudy (sparklestaraa):

:O i have never heard of that how do u do that log base thingy

OpenStudy (sparklestaraa):

maybe i should actually look at the lesson xD isn't that like trigonometry ?

OpenStudy (sparklestaraa):

@DanJS :( so whats it all mean